OpenStudy (anonymous):

Trigonometry help: Solve triangle ABC with A=51 degrees, b=21, and c=60. Capital letters are angle measures, lowercase letters are side lengths. Explanations would be helpful

OpenStudy (anonymous):

Assuming that this triangle is labeled in the traditional way (a is the side opposite angle A, and so on)… we can use the law of cosines to find the length of the third side. Is that what you mean by solve?

OpenStudy (anonymous):

I know i need to use the law of cosines, but for some reason when I use the formula the answer never comes out right

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

a^2 = b^2 + c^2 - 2 b c cos A Sound right?

OpenStudy (cwrw238):

a^2 = 21^2 + 60^2 - 2*21*60*cos 51 right?

OpenStudy (anonymous):

the answer comes down to about 30.9 when i plug the numbers in. But its a multiple choice question and the side length of a isnt that number

OpenStudy (anonymous):

I get 0.629 for cos(51 deg). Does that sound right?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

>>> a2 = 21**2 + 60**2 - 2*21*60*0.629 >>> a2 2455.92 >>> a2**0.5 49.55723963257034 Is that one of your answers?

OpenStudy (anonymous):

yes that is. i followed your steps and came to the same answer i think i might have been doing a mistake in the order of operations. thank you

OpenStudy (anonymous):

yw